A store owner has 12 pounds of pasta worth 70 cents a pound. She wants to mix it with pasta worth 45 cents a pound so that the total mixture can be sold for 55 cents a pound ( without any gain or loss). How much of the 45 cent pasta must she use?
Answer by checkley77(12844)

.70*12+.45x=.55(12+x)
8.4+.45x=6.6+.55x
.45x-.55x=6.6-8.4
.1x=1.8
x=1.8/.1
x=18 pounds of the $.45 pasta is needed.
Proof:
.70*12+.45*18=.55(12+18)
8.4+8.1=.55*30
16.5=16.5